### v1

# Quantum Mikheyev–Smirnov–Wolfenstein effect

Abstract
We study neutrino oscillations in background matter within the quantum field theory formalism where neutrino mass eigenstates are virtual particles. In this case, neutrino mass eigenstates are mixed owing to the interaction with matter. Assuming that neutrinos are Majorana particles, we find the exact propagators for massive neutrinos accounting for the interaction with matter by solving the analog of the Dyson equation. These propagators are used to calculate the transition probability which coincides with the prediction of the Mikheyev–Smirnov–Wolfenstein effect. Finally, we analyze the approximations made in our analysis.
1 Introduction
The experimental discoveries of neutrino oscillations (see, e.g., Ref. [1]) confirm that these particles are mixed and have nonzero masses. External fields are known to affect neutrino oscillations and even result in the new oscillations channels. It can be the case if neutrinos have magnetic moments and interact with external electromagnetic fields [2]. Then, besides flavor oscillations, one has the spin-flip, i.e. neutrino spin-flavor transitions are possible.
However, the external field influencing neutrino oscillations, more important from the point of view of possible applications, is the electroweak interaction with background matter. It can result in the significant amplification of the transition probability of oscillations known as the Mikheyev–Smirnov–Wolfenstein (MSW) effect [3, 4]. Nowadays, the MSW effect is the most plausible explanation of the solar neutrinos deficit [5].
Historically, neutrino flavor oscillations were described within the quantum mechanical approach. If one refers to neutrinos in vacuum, the neutrino mass eigenstates, which are the fundamental particles, propagate as plane waves. Recalculating the evolution for flavor eigenstates, one gets that the flavor content of a beam changes along the distance of the beam propagation. The quantum mechanical approach correctly describes neutrino oscillations in almost all reasonable situations.
Nevertheless, the quantum mechanical description has certain failures mainly from the methodological point of view. Some of these shortcomings are outlines in Refs. [6, 7]. Thus, a field theory approach for neutrino flavor oscillations is required. Numerous attempts to construct the theory of neutrino flavor oscillations have been made. Some of them are reviewed in Refs. [6, 7].
We mention one of these approaches based on the Quantum Field Theory (QFT), which was proposed in Refs. [8, 9]. It consists in the consideration of neutrinos, propagating from a source to a detector, as virtual particles. Some basic details are outlined shortly in Sec. 2. Then, this formalism was extensively studied, e.g., in Refs. [10, 11, 12, 13, 14].
Despite this QFT method links flavor oscillations and the foundations of QFT in a closest manner, it applies to either neutrino oscillations in vacuum or to the situation when an external field is diagonal in the mass eigenstate basis. It happens since one uses the propagators of neutrino mass eigenstates to compute the matrix element. In general situation, an external field mixes different mass eigenstates. Some exceptional cases when the interaction with external fields are diagonal in the mass basis are considered in Refs. [15, 16]. Thus, finding exact propagators is problematic even if one gets a propagator for a single mass eigenstate accounting for the interaction with an external field.
Neutrino oscillations in external fields were studied in a series of our works [15, 17, 18, 19], where we considered neutrinos as first quantized fields, i.e. we were based on the relativistic quantum mechanics. In that method, we used the solutions of wave equations exactly accounting for an external field. In Refs. [15, 17, 18, 19], we could reproduce the results for almost all reasonable oscillations channels. However, to get a result coinciding with the quantum mechanical approach, we had to take a quite broad initial neutrino wave packet.
In the present work, we study neutrino flavor oscillations in matter based on the QFT approach, i.e. we consider neutrino mass eigenstates as virtual particles. Despite of the difficulties mentioned above, we manage to account for the interaction of the mass eigenstates with matter exactly. Thus, based on reasonable approximations, we find the exact propagators of the mass eigenstates in matter. Eventually, we reproduce the MSW effect using our results.
This work is organized in the following way. We start in Sec. 2 with a brief description of the formalism for neutrino flavor oscillations within QFT. Then, in Sec. 3, we remind how mixed neutrinos interact with background matter. In Sec. 4, considering two flavor case, we derive the equations for the exact propagators of massive neutrinos in background matter. The developed formalism is adapted for Majorana neutrinos in Sec. 5. The exact propagators for massive Weyl neutrinos in matter are found in Sec. 6. These propagators are used in Sec. 7 to find the transition probability for neutrino flavor oscillations in matter within the QFT approach. Finally, we conclude in Sec. 8. The propagators of massive Weyl neutrinos accounting for the diagonal interaction with matter are derived in Appendix A. In Appendix B, we compute the specific 3D Fourier transforms of the propagators obtained in Appendix A.
2 Quantum field theory formalism for neutrino flavor oscillations
In this section, we outline the main features of the description of flavor oscillations where neutrinos are virtual particles. This formalism was proposed first in Ref. [8] and then independently studied in Ref. [9].
We suppose that a source and a detector of neutrinos are separated in space by the distance $\mathbf{L}$. The neutrino beam in a source is created by the interaction of charged leptons belonging to a certain flavor $\beta$ with the heavy nucleus $N$. The neutrino detection is implemented again by the interaction of charged leptons of the flavor $\alpha$ the a nucleus. If $\alpha\neq\beta$, one can speak about neutrino oscillations. The schematic Feynman diagram for the described process is depicted in Fig. 1.
Figure 1: The schematic illustration of neutrino oscillations in QFT.
The interaction between leptons and nuclei is described by the Lagrangian,
$\mathcal{L}_{\mathrm{int}}=\sqrt{2}G_{\mathrm{F}}j_{\mu}J^{\mu},$ (2.1)
where $G_{\mathrm{F}}$ is the coupling constant, e.g., it can be the Fermi constant,
$j_{\mu}=\sum_{\lambda}\bar{\nu}_{\lambda\mathrm{L}}\gamma_{\mu}l_{\lambda\mathrm{L}},$ (2.2)
is the leptonic current, $\gamma^{\mu}=(\gamma^{0},\bm{\gamma})$ are the Dirac matrices, $\nu_{\lambda}$ is the neutrino wavefunction corresponding to a certain flavor $\lambda=(e,\mu,\tau)$, $l_{\lambda}$ is the wavefunction of the charged lepton, and $J^{\mu}$ is the nuclear current. In Eq. (2.2), we assume that only left chiral projections of neutrinos and leptons participate in the interaction. For instance, $\nu_{\lambda\mathrm{L}}=P_{\mathrm{L}}\nu_{\lambda}$, where $P_{\mathrm{L}}=(1-\gamma^{5})/2$ and $\gamma^{5}=\mathrm{i}\gamma^{0}\gamma^{1}\gamma^{2}\gamma^{3}$ is the Dirac matrix.
Based on Eq. (2.1), one gets the $S$-matrix element for the process $(l_{\beta}N)\to\text{neutrinos}\to(l_{\alpha}N)$ in the form,
$S=-\frac{1}{2}\left(\sqrt{2}G_{F}\right)^{2}\int\mathrm{d}^{4}x\mathrm{d}^{4}y\left\langle l_{\alpha}\left|T\left\{j_{\mu}^{\dagger}(x)J^{\mu}(x)j^{\nu}(y)J_{\nu}^{\dagger}(y)\right\}\right|l_{\beta}\right\rangle,$ (2.3)
where $T$ stays for the chronological ordering of operators.
If charged leptons have the flavors $\beta$ in a source and $\alpha$ in a detector, then the corresponding operators in Eq. (2.2) act as
$\hat{l}_{\lambda}(y)\left|l_{\beta}\right\rangle=\frac{\delta_{\lambda\beta}}{\sqrt{2VE_{\beta}}}e^{-\mathrm{i}p_{\beta}y}u(p_{\beta})\left|0\right\rangle,\quad\left\langle l_{\alpha}\right|\hat{\bar{l}}_{\lambda}(x)=\frac{\delta_{\lambda\beta}}{\sqrt{2VE_{\alpha}}}e^{\mathrm{i}p_{\alpha}x}\bar{u}(p_{\alpha})\left\langle 0\right|$ (2.4)
where $V$ is the normalization volume, $p_{\alpha,\beta}^{\mu}=(E_{\alpha,\beta},\mathbf{p}_{\alpha,\beta})$ are the four-momenta of outgoing and incoming leptons, and $u(p_{\alpha,\beta})$ are their $c$-number wavefunctions. One implies in Eq. (2.4) that leptons propagate as plane waves.
The target nuclei are supposed to be quite heavy. Thus, their currents can be written in the form,
$J_{\mu}(x)\propto\delta_{\mu 0}\delta(\mathbf{x}-\mathbf{x}_{2}),\quad J_{\nu}(y)\propto\delta_{\nu 0}\delta(\mathbf{y}-\mathbf{x}_{1}),$ (2.5)
where $\mathbf{x}_{1,2}$ are the positions of the source and the detector, i.e. $\mathbf{L}=\mathbf{x}_{2}-\mathbf{x}_{1}$.
Flavor neutrinos $\nu_{\lambda}$ in Eq. (2.2) is the superposition of the mass eigenstates $\psi_{a}$,
$\nu_{\lambda}=\sum_{a}U_{\lambda a}\psi_{a},$ (2.6)
where $(U_{\lambda a})$ is the mixing matrix. If we consider only active neutrinos, $(U_{\lambda a})$ is a $3\times 3$ unitary matrix.
Using Eqs. (2.4)-(2.6), we rewrite Eq. (2.3) as
$S=$ $-\frac{G_{\mathrm{F}}^{2}e^{-\mathrm{i}\mathbf{p}_{\alpha}\mathbf{x}_{2}+\mathrm{i}\mathbf{p}_{\beta}\mathbf{x}_{1}}}{2V\sqrt{E_{\alpha}E_{\beta}}}\sum_{ab}U_{\alpha a}U_{\beta b}^{*}\int\mathrm{d}x_{0}\mathrm{d}y_{0}e^{\mathrm{i}E_{\alpha}x_{0}-\mathrm{i}E_{\beta}y_{0}}$
×$\times\bar{u}_{\mathrm{L}}(p_{\alpha})\gamma^{0}\left\langle 0\left|T\left\{\psi_{a\mathrm{L}}(x_{0},\mathbf{x}_{2})\bar{\psi}_{b\mathrm{L}}(y_{0},\mathbf{x}_{1})\right\}\right|0\right\rangle\gamma^{0}u_{\mathrm{L}}(p_{\beta}).$ (2.7)
Based on the Dirac matrices in the chiral representation, one transforms the factor in the integrand in Eq. (2.7),
$\bar{u}_{\mathrm{L}}(p_{\alpha})\gamma^{0}\left\langle 0\left|T\left\{\psi_{a\mathrm{L}}(x)\bar{\psi}_{b\mathrm{L}}(y)\right\}\right|0\right\rangle\gamma^{0}u_{\mathrm{L}}(p_{\beta})=\kappa_{-}^{\dagger}(p_{\alpha})\left\langle 0\left|T\left\{\eta_{a}(x)\eta_{b}^{\dagger}(y)\right\}\right|0\right\rangle\kappa_{-}(p_{\beta}),$ (2.8)
where $\eta_{a,b}$ and $\kappa(p_{\alpha,\beta})$ are the two component spinors defined by $\psi_{a,b\mathrm{L}}^{\mathrm{T}}=(0,\eta_{a,b})$ and $u_{\mathrm{L}}^{\mathrm{T}}(p_{\alpha,\beta})=(0,\kappa(p_{\alpha,\beta}))$. In Eq. (2.8), we assume that the charged leptons are ultrarelativistic, $E_{\alpha,\beta}\gg m_{\alpha,\beta}$. In this case, only $\kappa_{-}(p_{\alpha,\beta})=-\sqrt{2E_{\alpha,\beta}}w_{-}\neq 0$, where $w_{-}$ is the helicity amplitude given in Eq. (A.5).
If we consider massive Majorana neutrinos, we can represent them as Weyl fields $\eta_{a}$. In this situation, the expressions $\left\langle 0\left|T\left\{\eta_{a}(x)\eta_{b}^{\dagger}(y)\right\}\right|0\right\rangle=\Sigma_{ab}(x-y)$ are the propagators. We discuss them in Sec. 4.
The remaining integrals in Eq. (2.7) can be calculated using the new variables $t=x_{0}-y_{0}$ and $T=(x_{0}+y_{0})/2$. The final result has the form,
$S=-\frac{G_{\mathrm{F}}^{2}e^{-\mathrm{i}\mathbf{p}_{\alpha}\mathbf{x}_{2}+\mathrm{i}\mathbf{p}_{\beta}\mathbf{x}_{1}}}{2V\sqrt{E_{\alpha}E_{\beta}}}2\pi\delta(E_{\alpha}-E_{\beta})\mathcal{M}_{\beta\to\alpha},$ (2.9)
where the matrix element is
$\mathcal{M}_{\beta\to\alpha}=\sum_{ab}U_{\alpha a}U_{\beta b}^{*}\kappa_{-}^{\dagger}(p_{\alpha})\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\Sigma_{ab}(E,\mathbf{q})e^{\mathrm{i}\mathbf{qL}}\kappa_{-}(p_{\beta}).$ (2.10)
In Eq. (2.10), $\Sigma_{ab}(p)$ is the Fourier transform of $\Sigma_{ab}(x)$ and $E=(E_{\alpha}+E_{\beta})/2$.
The probability of the transition $\beta\to\alpha$ is $P_{\beta\to\alpha}\propto|\mathcal{M}_{\beta\to\alpha}|^{2}$. If virtual massive neutrinos propagate in vacuum, the propagators are diagonal, $\Sigma_{ab}\propto\delta_{ab}$. In this situation, the calculation of $P_{\beta\to\alpha}$ is quite straightforward; cf. Refs. [8, 9]. It gives one the expression for the probabilities of flavor neutrinos oscillations in vacuum. Our main goal is to find $\Sigma_{ab}$ and $\mathcal{M}_{\beta\to\alpha}$ for neutrinos propagating in background matter.
3 Mixed neutrinos in background matter
In this section, we briefly recall how neutrinos interact with background matter in frames of the field theory.
For simplicity, we consider the system of two mixed active flavor neutrinos $\nu_{\lambda}$, e.g., $\lambda=e,\mu$, in presence of background matter. This matter is supposed to be electroneutral, nonmoving, and unpolarized. The neutrino interaction with matter is treated in the forward scattering approximation. In this case, the Lagrangian for this system has the form,
$\mathcal{L}=\sum_{\lambda\lambda^{\prime}}\bar{\nu}_{\lambda}\left[\delta_{\lambda\lambda^{\prime}}\mathrm{i}\gamma^{\mu}\partial_{\mu}-m_{\lambda\lambda^{\prime}}-V_{\lambda\lambda^{\prime}}\frac{1}{2}\gamma^{0}(1-\gamma^{5})\right]\nu_{\lambda^{\prime}},$ (3.1)
where $(m_{\lambda\lambda^{\prime}})$ is the nondiagonal neutrino mass matrix. The matrix of the effective potentials for the neutrino interaction with matter $(V_{\lambda\lambda^{\prime}})$ is diagonal in the flavor basis in case of standard neutrino interactions, $V_{\lambda\lambda^{\prime}}=\delta_{\lambda\lambda^{\prime}}V_{\lambda}$. For example,
$V_{e}=\sqrt{2}G_{\mathrm{F}}\left(n_{e}-\frac{1}{2}n_{n}\right),\quad V_{\mu}=-\frac{G_{\mathrm{F}}}{\sqrt{2}}n_{n},$ (3.2)
for two neutrinos, $(\nu_{e},\nu_{\mu})$, in matter consisting of electrons, protons, and neutrons. In Eq. (3.2), $n_{e}=n_{p}$ and $n_{n}$ are the number densities of background fermions, and $G_{\mathrm{F}}=1.17\times 10^{-5}\,\text{GeV}^{-2}$ is the Fermi constant.
Then, we diagonalize the mass matrix and introduce the mass eigenstates, $\psi_{a}$, by means of the matrix transformation in Eq. (2.6), $\text{diag}(m_{1},m_{2})=U^{\dagger}(m_{\lambda\lambda^{\prime}})U$, where $m_{1,2}$ are the masses of $\psi_{1,2}$. In case of two neutrinos, $(U_{\lambda a})$ is represented in the form,
$(U_{\lambda a})=\left(\begin{array}[]{cc}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{array}\right),$ (3.3)
where $\theta$ is the vacuum mixing angle.
The Lagrangian in Eq. (3.1), rewritten in terms of the mass eigenstates, takes the form,
$\mathcal{L}=\sum_{ab}\bar{\psi}_{a}\left[\delta_{ab}\left(\mathrm{i}\gamma^{\mu}\partial_{\mu}-m_{a}\right)-g_{ab}\frac{1}{2}\gamma^{0}(1-\gamma^{5})\right]\psi_{b},$ (3.4)
where the matrix of effective potentials $(g_{ab})=U^{\dagger}VU$ is nondiagonal in general case. The Lagrangian in Eq. (3.4) is valid for both Dirac and Majorana neutrinos $\psi_{a}$. In Majorana case, we just replace $\frac{1}{2}\gamma^{0}(1-\gamma^{5})\to-\gamma^{0}\gamma^{5}$.
The wave equations for $\psi_{a}$, $a=1,2$, resulting from Eq. (3.4), have the form,
$\left[\mathrm{i}\gamma^{\mu}\partial_{\mu}-m_{1}-\frac{g_{1}}{2}\gamma^{0}(1-\gamma^{5})\right]\psi_{1}$ $=\frac{g}{2}\gamma^{0}(1-\gamma^{5})\psi_{2},$
$\left[\mathrm{i}\gamma^{\mu}\partial_{\mu}-m_{2}-\frac{g_{2}}{2}\gamma^{0}(1-\gamma^{5})\right]\psi_{2}$ $=\frac{g}{2}\gamma^{0}(1-\gamma^{5})\psi_{1},$ (3.5)
where $g\equiv g_{12}\neq 0$. One can see in Eq. (3.5) that the mass eigenstates are converted into each other in background matter.
4 Exact propagators of massive neutrinos in matter
In this section, we show how to obtain the exact propagators of neutrino mass eigenstates $\Sigma_{ab}$, which were defined in Sec. 2, in background matter. Thus, we consider a situation when the space between the source and the detector is filled with matter which a neutrino interacts with.
If a field $\psi$ obeys a wave equation $\mathcal{D}\psi=0$, where $\mathcal{D}$ is the differential operator, the Green function, e.g., a propagator $S$, can be found as the solution of the equation, $\mathcal{D}S=\delta(x)$, with the same differential operator. Bypassing the poles in the formal solution, one gets a Green function with required properties.
One can see in Eqs. (3.5) that it is problematic to solve the equation for a Green function even for two neutrinos. Indeed, one should consider the multiplet $\Psi=(\psi_{1},\psi_{2})$ and has to invert the $8\times 8$ matrix in Dirac indices while looking for a Green function. Nevertheless, we can tackle the problem of finding the propagators of mass eigenstates perturbatively and try to sum all the terms in a perturbation series.
We define the diagonal propagators of mass eigenstates as $S_{a}$, which obey the equations,
$\left[\mathrm{i}\gamma^{\mu}\partial_{\mu}-m_{a}-\frac{g_{a}}{2}\gamma^{0}(1-\gamma^{5})\right]S_{a}=\delta(x).$ (4.1)
These propagators exactly account for the diagonal interactions with matter $g_{a}$. In case of Dirac neutrinos, $S_{a}$ were found in Ref. [20] (see also Ref. [21]). The propagators in Eq. (4.1) for Majorana neutrinos are obtained in Appendix A.
To account for the transitions $1\leftrightarrow 2$ in matter, we notice that they are induced by the nondiagonal potential $G=\frac{g}{2}\gamma^{0}(1-\gamma^{5})$. For example, the exact propagators for the processes $1\to 1$ and $2\to 2$, $\Sigma_{11}$ and $\Sigma_{22}$, are the sums of the formal series,
$-\mathrm{i}\Sigma_{11}$ $=-\mathrm{i}S_{1}+(-\mathrm{i}S_{1})G(-\mathrm{i}S_{2})G(-\mathrm{i}S_{1})+\dotsb,$
$-\mathrm{i}\Sigma_{22}$ $=-\mathrm{i}S_{2}+(-\mathrm{i}S_{2})G(-\mathrm{i}S_{1})G(-\mathrm{i}S_{2})+\dotsb,$ (4.2)
whereas for the processes $1\to 2$ and $2\to 1$, $\Sigma_{12}$ and $\Sigma_{21}$, satisfy the expressions,
$-\mathrm{i}\Sigma_{12}$ $=(-\mathrm{i}S_{1})G(-\mathrm{i}S_{2})+(-\mathrm{i}S_{1})G(-\mathrm{i}S_{2})G(-\mathrm{i}S_{1})G(-\mathrm{i}S_{2})+\dotsb,$
$-\mathrm{i}\Sigma_{21}$ $=(-\mathrm{i}S_{2})G(-\mathrm{i}S_{1})+(-\mathrm{i}S_{2})G(-\mathrm{i}S_{1})G(-\mathrm{i}S_{2})G(-\mathrm{i}S_{1})+\dotsb.$ (4.3)
In Eqs. (4.2) and (4.3), $S_{1,2}=S_{1,2}(p)$ are the 4D Fourier images of the propagators $S_{1,2}(x)$. The same notice refers to $\Sigma_{ab}$. Note that the treatment of neutrino oscillations in matter analogous to that in Eqs. (4.2) and (4.3) was made in Ref. [22]. The analytical expressions in Eqs. (4.2) and (4.3) are illustrated in Fig. 2.
Figure 2: Feynman diagrams corresponding to Eqs. (4.2) and (4.3). 4D Fourier images of the bare propagators $-\mathrm{i}S_{a}$ are depicted by thin solid lines, whereas thick solid lines correspond to the dressed propagators $-\mathrm{i}\Sigma_{ab}$ [23].
The sum of the series in Eqs. (4.2) and (4.3) can be represented in the form of the Dyson equation for an exact Green function,
$(-\mathrm{i}S_{1})^{-1}$ $=(-\mathrm{i}\Sigma_{11})^{-1}+G(-\mathrm{i}S_{2})G,$
$(-\mathrm{i}S_{2})^{-1}$ $=(-\mathrm{i}\Sigma_{22})^{-1}+G(-\mathrm{i}S_{1})G,$ (4.4)
and
$\left[(-\mathrm{i}S_{1})G(-\mathrm{i}S_{2})\right]^{-1}$ $=(-\mathrm{i}\Sigma_{12})^{-1}+G,$
$\left[(-\mathrm{i}S_{2})G(-\mathrm{i}S_{1})\right]^{-1}$ $=(-\mathrm{i}\Sigma_{21})^{-1}+G.$ (4.5)
The solutions of Eq. (4.4) is quite nontrivial especially for Dirac neutrinos.
One can notice that Eq. (4.5) implies the inversion of the effective potential $G$. However, it contains the projection $(1-\gamma^{5})/2$ for Dirac neutrinos. It means that $G^{-1}$ does not exist. Nevertheless, the developed formalism is well defined for Majorana particles. This result is in agreement with the statement (see, e.g., Refs. [24, 25]) that neutrino mass eigenstates are likely to be Majorana particles.
5 Reformulation of the formalism for Majorana mass eigenstates
In this section, we rewrite the basic expressions in Secs.3 and 4 for Majorana neutrinos.
Since we found in Sec. 4 that transitions between mass eigenstates in matter can be described for Majorana neutrinos, we should rewrite the basic expressions for this kind of particles. Instead of dealing with a four components wave function, we use the Weyl representation for a Majorana neutrino in terms of the two component spinor $\eta$. It is made in order to avoid the constraint, i.e. the Majorana condition, in the quantization of the fields. If we use the chiral representation of Dirac matrices, the four component Majorana wave function has the form $\psi^{\mathrm{T}}=(\mathrm{i}\sigma_{2}\eta^{*},\eta)$, where $\eta$ corresponds to a left-handed Weyl neutrino, and $\sigma_{2}$ is the Pauli matrix.
The interaction of a Weyl neutrino mass eigenstate with matter takes the form,
$\mathcal{L}_{\mathrm{int}}=\sum_{ab}g_{ab}\bar{\psi}_{a\mathrm{L}}\gamma^{0}\gamma^{5}\psi_{b\mathrm{L}}=-\sum_{ab}g_{ab}\eta_{a}^{\dagger}\eta_{b}.$ (5.1)
The analog of Eq. (3.5) for Weyl neutrinos is (see also Appendix A)
$\mathrm{i}\dot{\eta}_{1}+[(\bm{\sigma}\mathbf{p})-g_{1}]\eta_{1}+\mathrm{i}m_{1}\sigma_{2}\eta_{1}^{*}$ $=g\eta_{2},$
$\mathrm{i}\dot{\eta}_{2}+[(\bm{\sigma}\mathbf{p})-g_{2}]\eta_{2}+\mathrm{i}m_{2}\sigma_{2}\eta_{2}^{*}$ $=g\eta_{1},$ (5.2)
where $\bm{\sigma}=(\sigma_{1},\sigma_{2},\sigma_{3})$ are the Pauli matrices, $\mathbf{p}=-\mathrm{i}\nabla$, and a dot means the time derivative.
The relation between $(g_{1,2},g)$ and $V_{e,\mu}$ is the same as for Dirac neutrinos. The main results of Sec. 4, e.g., Eqs. (4.4) and (4.5), remain valid for Weyl neutrinos as well. However, $G=g$ is the scalar quantity.
[…]
6 Exact propagators of massive Weyl neutrinos
In this section, we solve Eqs. 4.4 and 4.5 for Majorana neutrinos.
We represent the 4D Fourier image of the propagator $S_{a}$ in Eq. (A.10) in the form,
$S_{a}(p)=\mathcal{A}_{a}(p)+(\bm{\sigma}\hat{p})\mathcal{B}_{a}(p),$ (6.1)
where both $\mathcal{A}_{a}$ and $\mathcal{B}_{a}$ contain four terms. Then, after tedious but straightforward calculations, one finds the solution of Eqs. (4.4) and (4.5),
$\Sigma_{11}$ $=\frac{\mathcal{A}_{1}+(\bm{\sigma}\hat{p})\mathcal{B}_{1}+g^{2}(\mathcal{A}_{1}^{2}-\mathcal{B}_{1}^{2})(\mathcal{A}_{2}-(\bm{\sigma}\hat{p})\mathcal{B}_{2})}{(1+g^{2}\mathcal{A}_{1}\mathcal{A}_{2})^{2}+(1+g^{2}\mathcal{B}_{1}\mathcal{B}_{2})^{2}-1-g^{4}(\mathcal{A}_{1}^{2}\mathcal{B}_{2}^{2}+\mathcal{A}_{2}^{2}\mathcal{B}_{1}^{2})},$
$\Sigma_{22}$ $=\frac{\mathcal{A}_{2}+(\bm{\sigma}\hat{p})\mathcal{B}_{2}+g^{2}(\mathcal{A}_{2}^{2}-\mathcal{B}_{2}^{2})(\mathcal{A}_{1}-(\bm{\sigma}\hat{p})\mathcal{B}_{1})}{(1+g^{2}\mathcal{A}_{1}A_{2})^{2}+(1+g^{2}B_{1}B_{2})^{2}-1-g^{4}(\mathcal{A}_{1}^{2}B_{2}^{2}+\mathcal{A}_{2}^{2}B_{1}^{2})},$
$\Sigma_{12}$ $=\Sigma_{21}=\frac{\mathrm{i}g\left[\mathcal{A}_{1}\mathcal{A}_{2}+\mathcal{B}_{1}\mathcal{B}_{2}-g^{2}(\mathcal{A}_{1}^{2}-\mathcal{B}_{1}^{2})(\mathcal{A}_{2}^{2}-\mathcal{B}_{2}^{2})-(\bm{\sigma}\hat{p})(\mathcal{A}_{1}\mathcal{B}_{2}-\mathcal{B}_{1}\mathcal{A}_{2})\right]}{(1+g^{2}\mathcal{A}_{1}\mathcal{A}_{2})^{2}+(1+g^{2}\mathcal{B}_{1}\mathcal{B}_{2})^{2}-1-g^{4}(\mathcal{A}_{1}^{2}\mathcal{B}_{2}^{2}+\mathcal{A}_{2}^{2}\mathcal{B}_{1}^{2})},$ (6.2)
which is valid for neutrinos with arbitrary energies.
Further analysis of Eq. (6.2) is quite difficult since $\mathcal{A}_{a}$ and $\mathcal{B}_{a}$ have the complicated structure. However, as shown in Appendix B, the main contribution to the matrix element involving the 3D Fourier transform in Eq. (B.1) comes from $I_{1}$ and $J_{1}$ terms. That is why, we keep only one term in $\mathcal{A}_{a}$ and $\mathcal{B}_{a}$, i.e.
$\mathcal{A}_{a}=-\mathcal{B}_{a}=\frac{\mathrm{i}\lambda_{a-}^{2}}{2(p_{0}-E_{a-}+\mathrm{i}0)}.$ (6.3)
Note that Eq. (6.3) is valid for ultrarelativistic neutrinos.
To proceed, we assume that the interaction with matter is weak. In this approximation, we keep the terms up to $\sim g^{2}$ in Eq. (6.2). Using Eq. (6.3), we rewrite Eq. (6.2) in the form,
$\Sigma_{11}(p_{0},\mathbf{p})$ $=\frac{\mathrm{i}[1-(\bm{\sigma}\hat{p})](p_{0}-E_{2-})}{2[(p_{0}-E_{1-}+\mathrm{i}0)(p_{0}-E_{2-}+\mathrm{i}0)-g^{2}]},$
$\Sigma_{22}(p_{0},\mathbf{p})$ $=\frac{\mathrm{i}[1-(\bm{\sigma}\hat{p})](p_{0}-E_{1-})}{2[(p_{0}-E_{1-}+\mathrm{i}0)(p_{0}-E_{2-}+\mathrm{i}0)-g^{2}]},$
$\Sigma_{12}(p_{0},\mathbf{p})$ $=\frac{\mathrm{i}g[1-(\bm{\sigma}\hat{p})]}{2[(p_{0}-E_{1-}+\mathrm{i}0)(p_{0}-E_{2-}+\mathrm{i}0)-g^{2}]}.$ (6.4)
The exact propagators $\Sigma_{ab}$, accounting for the interaction of massive neutrinos with matter, are used in Eq. (2.10) to calculate the matrix element.
7 Transition probability for oscillations in matter
In this section, we apply our results to derive the transition probability of neutrino oscillations in matter basing on QFT.
Let us consider the particular oscillations channel, $e\to\mu$. Using Eqs. (2.10) and (3.3), we get that the matrix element is
$\mathcal{M}_{e\to\mu}\propto\kappa_{-}^{\dagger}(p_{\mu})\left[\sin 2\theta\frac{1}{2}(\Sigma_{22}-\Sigma_{11})+\cos 2\theta\Sigma_{12}\right]\kappa_{-}(p_{e}),$ (7.1)
where the exact propagators $\Sigma_{ab}(p_{0},\mathbf{p})$ are given in Eq. (6.4). To get the final form of the matrix element in Eq. (7.1) one should put $p_{0}=E$, calculate the 3D Fourier transform, as well as average the result over the incoming and outgoing left-handed charged leptons.
We make the above calculations in details, e.g., for the $\Sigma_{12}$ contribution. After substituting $p_{0}\to E$, we transform the factor in the integrand
$\frac{1}{(E-E_{1-}+\mathrm{i}0)(E-E_{2-}+\mathrm{i}0)-g^{2}}=\frac{1}{\mathcal{E}_{+}-\mathcal{E}_{-}}\left(\frac{1}{E-\mathcal{E}_{+}+\mathrm{i}0}-\frac{1}{E-\mathcal{E}_{-}+\mathrm{i}0}\right),$ (7.2)
where
±$\mathcal{E}_{\pm}=\frac{E_{1-}+E_{2-}}{2}\pm\sqrt{\left(\frac{E_{1-}-E_{2-}}{2}\right)^{2}+g^{2}}.$ (7.3)
Note that $\mathcal{E}_{\pm}$ depend quite non-trivially on $q$.
Based on Eqs. (7.2) and (7.3), we compute, e.g., the integral
$I=\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{e^{\mathrm{i}\mathbf{qL}}}{(\mathcal{E}_{+}-\mathcal{E}_{-})(E-\mathcal{E}_{+}+\mathrm{i}0)},$ (7.4)
which contributes to the 3D Fourier transform of $\Sigma_{12}$. Analogously to Appendix B, we use the cylindrical coordinates $\mathbf{q}=(\rho,z,\phi)$ and assume that $\mathbf{L}$ is along the positive direction of $\mathbf{e}_{z}$. Then,
$I=-\frac{\mathrm{i}}{2\pi}\int_{0}^{+\infty}\rho\mathrm{d}\rho\sum_{\left\{z_{+}\right\}}\left.\frac{e^{\mathrm{i}zL}}{(\mathcal{E}_{+}-\mathcal{E}_{-})\frac{\mathrm{d}\mathcal{E}_{+}}{\mathrm{d}z}}\right|_{z=z_{+}},$ (7.5)
where $\left\{z_{+}(\rho)\right\}$ are the roots of the equation $E=\mathcal{E}_{+}$. Since $L>0$, we close the contour in the upper half-plane. All roots $\left\{z_{+}\right\}$ are in the upper half-plane as well.
In general case, the equation $E=\mathcal{E}_{+}$ is equivalent to the algebraic equation of the 8th degree. However, if neutrinos are ultrarelativistic, i.e. $E\gg m_{1,2}$, we can solve it perturbatively. First, we represent Eq. (7.3) in the form,
$\mathcal{E}_{+}\approx q+\sqrt{\left(\frac{\Delta m^{2}}{4q}+\frac{g_{2}-g_{1}}{2}\right)^{2}+g^{2}},$ (7.6)
where $\Delta m^{2}=m_{2}^{2}-m_{1}^{2}$. In the first approximation, we take $q(z_{+})=\sqrt{z_{+}^{2}+\rho^{2}}\approx\mathcal{E}_{+}(E)$, where
$\mathcal{E}_{+}(E)=E+\sqrt{\left(\frac{\Delta m^{2}}{4E}+\frac{g_{2}-g_{1}}{2}\right)^{2}+g^{2}}.$ (7.7)
[…]
$I=-\frac{Ee^{\mathrm{i}\mathcal{E}_{+}(E)L}}{2\pi L\sqrt{\left(\frac{\Delta m^{2}}{4E}+\frac{g_{2}-g_{1}}{2}\right)^{2}+g^{2}}}.$ (7.8)
The rest of the integrals contributing to the matrix element in Eq. (7.1) is computed analogously.
We present the final result for the matrix element in Eq. (7.1),
$\mathcal{M}_{e\to\mu}=$ $-\frac{\left(\frac{\Delta m^{2}}{4E}+\frac{g_{2}-g_{1}}{2}\right)\sin 2\theta+g\cos 2\theta}{\sqrt{\left(\frac{\Delta m^{2}}{4E}+\frac{g_{2}-g_{1}}{2}\right)^{2}+g^{2}}}\sin\left(\sqrt{\left(\frac{\Delta m^{2}}{4E}+\frac{g_{2}-g_{1}}{2}\right)^{2}+g^{2}|}L|\right)$
×$\times\frac{\mathrm{i}Ee^{\mathrm{i}E|L|}}{2\pi|L|}\kappa_{-}^{\dagger}(p_{\mu})\left[1-(\bm{\sigma}\hat{L})\right]\kappa_{-}(p_{e}).$ (7.9)
Analogously to Appendix B, we assume that the neutrino propagation distance is large in deriving of Eq. (7.9). If the incoming and outgoing leptons are ultrarelativistic and their momenta are almost along the same direction, the quantity $\kappa_{-}^{\dagger}(p_{\mu})\left[1-(\bm{\sigma}\hat{L})\right]\kappa_{-}(p_{e})$ is just a nonzero constant factor.
The effective potentials of the mass eigenstates interaction with matter are
$g_{1}=V_{e}\cos^{2}\theta+V_{\mu}\sin^{2}\theta,\quad g_{1}=V_{e}\sin^{2}\theta+V_{\mu}\cos^{2}\theta,\quad g=(V_{e}-V_{\mu})\sin\theta\cos\theta,$ (7.10)
where $V_{e,\mu}$ are given in Eq. (3.2). Finally, we get that the probability of the process $e\to\mu$ contains the factor,
$P_{e\to\mu}\propto$ $\frac{\left(\frac{\Delta m^{2}}{4E}\sin 2\theta\right)^{2}}{\left(\frac{\Delta m^{2}}{4E}\sin 2\theta\right)^{2}+\left(\frac{\Delta m^{2}}{4E}\cos 2\theta-\frac{G_{\mathrm{F}}n_{e}}{\sqrt{2}}\right)^{2}}$
×$\times\sin^{2}\left(\sqrt{\left(\frac{\Delta m^{2}}{4E}\sin 2\theta\right)^{2}+\left(\frac{\Delta m^{2}}{4E}\cos 2\theta-\frac{G_{\mathrm{F}}n_{e}}{\sqrt{2}}\right)^{2}}|L|\right),$ (7.11)
which coincides with the prediction of the MSW effect.
8 Conclusion
In the present work, we have studied neutrino flavor oscillations in background matter in frames of the QFT approach where massive neutrinos are virtual particles. We have started in Sec. 2 with reminding of the basic concepts of this method. The calculation of the matrix element for the process corresponding to oscillating neutrinos involves propagators of mass eigenstates. We have established in Sec. 3 that mass eigenstates are mixed in background matter in general case. It was the main difficulty in the application of QFT for neutrino oscillations in matter.
In Sec. 4, we have derived the equations for the exact propagators, which account for the neutrino interaction with matter. These equations are analogous to the Dyson equations. Moreover, we have established that the developed formalism is not applicable for the Dirac mass eigenstates. For this reason, in Sec. 5, we have reformulated the basic concepts of our approach in terms of Weyl spinors, which are the representation of Majorana neutrinos. The exact propagators for this kind of neutrinos have been found in Sec. 6.
In Sec. 7, we have applied of findings to a particular case of $\nu_{e}\to\nu_{\mu}$ oscillations. We have derived the matrix element and the corresponding transition probability. One can see in Eq. (7.11) that it reproduces the MSW effect.
The diagonal propagators of Weyl neutrinos have been found in Appendix A. The specific 3D Fourier transforms of the propagators have been computed in Appendix B.
Despite it is difficult to get something beyond Eq. (7.11) which was originally derived within the quantum mechanical approach, it is instructive to apply the exact formalism of QFT for neutrino flavor oscillations mainly from the methodological point of view. It allows one to analyze the validity of approximations made to obtain Eq. (7.11). The main assumptions used in our work are (i) that neutrinos are ultrarelativistic particles; and (ii) the weakness of the matter interaction.
The assumption (i) was used first in Sec. 5, where it allowed us to exclude the involvement of antineutrinos in the series in Eqs. (4.2) and (4.3). Without this assumption, the summation of the diagrams in Fig. 1 would be quite nontrivial. Second, the assumption (i) was utilized in Appendix B in the computation of the integrals $I_{i}$ and $J_{i}$. It allowed us to show that the contribution of $I_{1}$ and $J_{1}$ is the leading one. Thus, we could reduce the number of terms in $S_{a}$; cf. Eq. (6.3). Third, the assumption (i) was used in the calculation of the 3D Fourier transform in Eqs. (7.6)-(7.8) resulting in the matrix element in Eq. (7.9). Thus, we can see that the ultrarelativity of neutrinos is crucial for Eq. (7.11).
The assumption (ii) was used to simplify the propagators and jump from Eq. (6.2) to Eq. (6.4). In fact, it allowed us to get the $g^{2}$-term under the square root in Eq. (7.3). In principle, one can study corrections to Eq. (7.11) for neutrinos propagating in extremely dense matter.
Appendix A Massive Weyl neutrino propagators in matter
In this Appendix, we derive the propagators of massive Weyl neutrinos in matter. The Lagrangian for a single mass eigenstate, described by the two component wave function $\eta_{a}$, having the mass $m_{a}$ and interacting with nonmoving and unpolarized matter is
$\mathcal{L}=\mathrm{i}\eta_{a}^{\dagger}(\sigma^{\mu}\partial_{\mu})\eta_{a}-g\eta_{a}^{\dagger}\eta_{a}+\frac{\mathrm{i}}{2}m_{a}\eta_{a}^{\dagger}\sigma_{2}\eta_{a}^{*}-\frac{\mathrm{i}}{2}m\eta_{a}^{\mathrm{T}}\sigma_{2}\eta_{a},$ (A.1)
where $\sigma^{\mu}=(1,-\bm{\sigma})$, and $g_{a}$ is given in Sec. 3. Based on Eq. (A.1), we obtain the wave equation for $\eta_{a}$
$\mathrm{i}\dot{\eta}_{a}+[(\bm{\sigma}\mathbf{p})-g_{a}]\eta_{a}+\mathrm{i}m_{a}\sigma_{2}\eta_{a}^{*}=0.$ (A.2)
Analogous equation is obtained for $\eta_{a}^{*}$.
The general solution of Eq. (A.2) is
$\eta_{a}(\mathbf{x},t)=$ $\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3/2}}\bigg{[}\lambda_{a-}\left(a_{-}w_{-}e^{-\mathrm{i}E_{a-}t+i\mathbf{px}}+A_{a-}a_{-}^{\dagger}w_{+}e^{\mathrm{i}E_{a-}t-\mathrm{i}\mathbf{px}}\right)$
$+\lambda_{a+}\left(A_{a+}a_{+}w_{+}e^{-\mathrm{i}E_{a+}t+\mathrm{i}\mathbf{px}}+a_{+}^{\dagger}w_{-}e^{\mathrm{i}E_{a+}t-\mathrm{i}\mathbf{px}}\right)\bigg{]},$ (A.3)
where $E_{a\pm}=\sqrt{m_{a}^{2}+(p\mp g_{a})^{2}},$ are the energies for right and left polarized neutrinos, $a_{\pm}^{\dagger}$ and $a_{\pm}$ are the creation and annihilation operators for certain helicities,
$\lambda_{a\pm}^{2}=\frac{E_{a\pm}+p\mp g_{a}}{2E_{a\pm}},\quad A_{a\pm}=\mp\frac{m_{a}}{E_{a\pm}+p\mp g_{a}},$ (A.4)
and
$w_{+}=\left(\begin{array}[]{c}e^{-\mathrm{i}\phi/2}\cos\vartheta/2\\ e^{\mathrm{i}\phi/2}\sin\vartheta/2\end{array}\right),\quad w_{-}=\left(\begin{array}[]{c}-e^{-\mathrm{i}\phi/2}\sin\vartheta/2\\ e^{\mathrm{i}\phi/2}\cos\vartheta/2\end{array}\right),$ (A.5)
are the helicity amplitudes. In Eq. (A.5), the angles $\vartheta$ and $\phi$ fix the direction of the neutrino momentum $\mathbf{p}$. We mention some important properties of $w_{\pm}$,
$w_{\pm}\otimes w_{\pm}^{\dagger}=\frac{1}{2}\left[1\pm\bm{\sigma}\hat{p}\right],\quad w_{\pm}\otimes w_{\mp}^{\mathrm{T}}=\pm\frac{1}{2}\left(1\pm\bm{\sigma}\hat{p}\right)\mathrm{i}\sigma_{2},$ (A.6)
which can be checked by means direct calculations using Eq. (A.5).
Note that $\eta_{a}(\mathbf{x},t)$ in Eq. (A.3) and analogous $\eta_{a}^{*}(\mathbf{x},t)$ obey the canonical equal time anticommutation relation $\left\{\eta_{a}(\mathbf{x},t),\eta_{a}^{*}(\mathbf{y},t)\right\}=\delta(\mathbf{x}-\mathbf{y})$ provided that $\left\{a_{\pm}(\mathbf{p}),a_{\pm}^{\dagger}(\mathbf{q})\right\}=\delta(\mathbf{p}-\mathbf{q})$ and all other anticommutators vanish. Thus, we have constructed the canonical quantization of Weyl neutrinos in background matter.
We define two propagators for Weyl neutrinos
$S_{a}(x-y)=$ $\left\langle 0\left|T[\eta_{a}(x)\eta_{a}^{\dagger}(y)]\right|0\right\rangle=\theta(x_{0}-y_{0})\left\langle 0\left|\eta_{a}(x)\eta_{a}^{\dagger}(y)\right|0\right\rangle$
$-\theta(y_{0}-x_{0})\left\langle 0\left|\eta_{a}^{*}(y)\eta_{a}^{\mathrm{T}}(x)\right|0\right\rangle,$ (A.7)
and
$\tilde{S}_{a}(x-y)=$ $\left\langle 0\left|T[\eta_{a}(x)\eta_{a}^{\mathrm{T}}(y)]\right|0\right\rangle=\theta(x_{0}-y_{0})\left\langle 0\left|\eta_{a}(x)\eta_{a}^{\mathrm{T}}(y)\right|0\right\rangle$
$-\theta(y_{0}-x_{0})\left\langle 0\left|\eta_{a}(y)\eta_{a}^{\mathrm{T}}(x)\right|0\right\rangle,$ (A.8)
where $\theta(t)$ is the Heaviside step function. Using Eqs. (A.3), (A.4), and (A.6), as well as the identity
$e^{\pm\mathrm{i}Et}\theta(\mp t)=\pm\frac{1}{2\pi\mathrm{i}}\int_{-\infty}^{+\infty}\frac{e^{-\mathrm{i}p_{0}t}\mathrm{d}p_{0}}{p_{0}\pm E\mp\mathrm{i}0},$ (A.9)
[…]
$S_{a}(x)=$ $\frac{\mathrm{i}}{2}\int\frac{\mathrm{d}^{4}p}{(2\pi)^{4}}e^{-\mathrm{i}px}\bigg{\{}\bigg{[}\lambda_{a-}^{2}\left(\frac{1}{p_{0}-E_{a-}+\mathrm{i}0}+\frac{A_{a-}^{2}}{p_{0}+E_{a-}-\mathrm{i}0}\right)$
$+$ $\lambda_{a+}^{2}\left(\frac{1}{p_{0}+E_{a+}-\mathrm{i}0}+\frac{A_{a+}^{2}}{p_{0}-E_{a+}+\mathrm{i}0}\right)\bigg{]}$
$-(\bm{\sigma}\hat{p})\bigg{[}\lambda_{a-}^{2}\left(\frac{1}{p_{0}-E_{a-}+\mathrm{i}0}+\frac{A_{a-}^{2}}{p_{0}+E_{a-}-\mathrm{i}0}\right)$
-- $\lambda_{a+}^{2}\left(\frac{1}{p_{0}+E_{a+}-\mathrm{i}0}+\frac{A_{a+}^{2}}{p_{0}-E_{a+}+\mathrm{i}0}\right)\bigg{]}\bigg{\}},$ (A.10)
$\tilde{S}_{a}(x)=$ $\frac{m\sigma_{2}}{2}\int\frac{\mathrm{d}^{4}p}{(2\pi)^{4}}e^{-\mathrm{i}px}\left[\frac{\left(1-\bm{\sigma}\hat{p}\right)}{p_{0}^{2}-E_{a-}^{2}+\mathrm{i}0}+\frac{\left(1+\bm{\sigma}\hat{p}\right)}{p_{0}^{2}-E_{a+}^{2}+\mathrm{i}0}\right].$ (A.11)
Note that, if one turns off the neutrino interaction with matter, then $E_{a+}=E_{a-}=\sqrt{p^{2}+m_{a}^{2}}$. Thus, $S_{a}$ and $\tilde{S}_{a}$ in Eqs. (A.10) and (A.11) reproduce the known vacuum propagators for a massive Weyl neutrino in Ref. [26].
Appendix B 3D Fourier transform of propagators
In this Appendix, we calculate the Fourier transform of the propagator of a single massive Weyl neutrino in matter. We use the corresponding propagators in Appendix A.
In Sec. 2, we obtain that the matrix element contains the Fourier transform of the propagator in the form,
$\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}S_{a}(E,\mathbf{q})e^{\mathrm{i}\mathbf{qL}},$ (B.1)
where $E$ is the mean energy of incoming and outgoing leptons and $\mathbf{L}$ is the vector between the source and the detector. One can see in Eq. (A.10) that the Fourier image of the propagator contains eight terms. We provide the detailed calculations only for two of them,
$I_{1}=\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a-}^{2}e^{\mathrm{i}\mathbf{qL}}}{E-E_{a-}+\mathrm{i}0},\quad J_{1}=\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a-}^{2}(\bm{\sigma}\hat{q})e^{\mathrm{i}\mathbf{qL}}}{E-E_{a-}+\mathrm{i}0},$ (B.2)
The rest of the integrals is computed analogously.
We use the cylindrical coordinates $\mathbf{q}=(\rho,z,\phi)$ and suppose that $\mathbf{L}\parallel\mathbf{e}_{z}$. For example,
$I$ =${}_{1}=\frac{1}{(2\pi)^{2}}\int_{0}^{+\infty}\rho\mathrm{d}\rho\int_{-\infty}^{+\infty}\mathrm{d}z\frac{(E_{a-}+q+g_{a})e^{\mathrm{i}zL}}{2E_{a-}(E-\sqrt{m_{a}^{2}+(\sqrt{z^{2}+\rho^{2}}+g_{a})^{2}}+\mathrm{i}0)}.$ (B.3)
In the integrand in Eq. (B.3), we take into account that $E_{a-}=\sqrt{m_{a}^{2}+(\sqrt{z^{2}+\rho^{2}}+g_{a})^{2}}$ and $q=\sqrt{z^{2}+\rho^{2}}$.
We expand the function $Z(z)=E-\sqrt{m_{a}^{2}+(\sqrt{z^{2}+\rho^{2}}+g_{a})^{2}}$ in series near the pole $z_{0}$, $Z(z_{0})=0$,
$Z\approx Z^{\prime}(z-z_{0})+\dotsc,\quad Z^{\prime}=\frac{\mathrm{d}}{\mathrm{d}z}\left[E-\sqrt{m_{a}^{2}+(\sqrt{z^{2}+\rho^{2}}+g_{a})^{2}}\right]_{z=z_{0}}=-\frac{z_{0}(q_{0}+g_{a})}{q_{0}E},$ (B.4)
[…]
$I_{1}=$ $-\frac{\mathrm{i}\rho_{0}(E+\rho_{0}+g_{a})}{4\pi(\rho_{0}+g_{a})}\left(\int_{0}^{\rho_{0}}\rho\mathrm{d}\rho\frac{e^{\mathrm{i}\sqrt{\rho_{0}^{2}-\rho^{2}}L}}{\sqrt{\rho_{0}^{2}-\rho^{2}}}-\mathrm{i}\int_{\rho_{0}}^{\infty}\rho\mathrm{d}\rho\frac{e^{-\sqrt{\rho^{2}-\rho_{0}^{2}}L}}{\sqrt{\rho^{2}-\rho_{0}^{2}}}\right)$
$=-\frac{\left(E+\sqrt{E^{2}-m_{a}^{2}}\right)\left(\sqrt{E^{2}-m_{a}^{2}}-g\right)}{4\pi L\sqrt{E^{2}-m_{a}^{2}}}e^{\mathrm{i}(\sqrt{E^{2}-m_{a}^{2}}-g_{a})L}.$ (B.5)
[…]
$I_{1}$ $=-\frac{\left(E+\sqrt{E^{2}-m_{a}^{2}}\right)\left(\sqrt{E^{2}-m_{a}^{2}}-g_{a}\right)}{4\pi|L|\sqrt{E^{2}-m_{a}^{2}}}e^{\mathrm{i}(\sqrt{E^{2}-m_{a}^{2}}-g_{a})|L|},$ (B.6)
where we incorporate both positive and negative values of $L$.
The integrand in $J_{1}$ in Eq. (B.2) contains the matrix factor
$(\bm{\sigma}\hat{q})=\frac{1}{\sqrt{z^{2}+\rho^{2}}}\left(\begin{array}[]{cc}z&\rho e^{-\mathrm{i}\phi}\\ \rho e^{\mathrm{i}\phi}&-z\end{array}\right),$ (B.7)
in addition to that in Eq. (B.3). The integration over $\phi$, analogous to that in Eq. (B.3), removes the off-diagonal terms in the final result. Thus, we should calculate only the diagonal elements which differ from these in Eq. (B.3) by the additional factors ±$\pm z/\sqrt{z^{2}+\rho^{2}}$ in the integrands. Since the computation is almost identical to that for $I_{1}$, we just present the final result,
$\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a-}^{2}e^{\mathrm{i}\mathbf{qL}}}{E-E_{a-}+\mathrm{i}0}\left(\pm\frac{z}{\sqrt{z^{2}+\rho^{2}}}\right)=\mp\frac{\left(E+\sqrt{E^{2}-m_{a}^{2}}\right)\left(\sqrt{E^{2}-m_{a}^{2}}-g_{a}\right)}{4\pi L\sqrt{E^{2}-m_{a}^{2}}}\\ \times e^{\mathrm{i}(\sqrt{E^{2}-m_{a}^{2}}-g_{a})|L|}\left(1+\frac{\mathrm{i}}{|L|(\sqrt{E^{2}-m_{a}^{2}}-g_{a})}\right),$ (B.8)
which is valid for arbitrary values of $L$. Taking into account the chosen direction of $\mathbf{L}$, the final result for $J_{1}$ reads
$J_{1}=$ $\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a-}^{2}(\bm{\sigma}\hat{q})e^{\mathrm{i}\mathbf{qL}}}{E-E_{a-}+\mathrm{i}0}=-(\bm{\sigma}\hat{L})\frac{\left(E+\sqrt{E^{2}-m_{a}^{2}}\right)\left(\sqrt{E^{2}-m_{a}^{2}}-g_{a}\right)}{4\pi|L|\sqrt{E^{2}-m_{a}^{2}}}$
×$\times e^{\mathrm{i}(\sqrt{E^{2}-m_{a}^{2}}-g_{a})|L|}\left(1+\frac{\mathrm{i}}{|L|(\sqrt{E^{2}-m_{a}^{2}}-g_{a})}\right).$ (B.9)
Considering the limit of the great distance of the neutrino propagation, one has that $J_{1}\approx(\bm{\sigma}\hat{L})I_{1}$.
Eventually, we list the results for the rest of the terms in Eq. (A.10). They are
$I_{2}$ $=\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a+}^{2}A_{a+}^{2}e^{\mathrm{i}\mathbf{qL}}}{E-E_{a+}+\mathrm{i}0}=-\frac{m_{a}^{2}\left(\sqrt{E^{2}-m_{a}^{2}}+g_{a}\right)}{4\pi|L|\left(E+\sqrt{E^{2}-m_{a}^{2}}\right)\sqrt{E^{2}-m_{a}^{2}}}$
×$\times e^{\mathrm{i}(\sqrt{E^{2}-m_{a}^{2}}+g_{a})|L|},$
$I_{3}$ $=\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a-}^{2}A_{a-}^{2}e^{\mathrm{i}\mathbf{qL}}}{E+E_{a-}-\mathrm{i}0}=-\frac{m_{a}^{2}\left(\sqrt{E^{2}-m_{a}^{2}}-g_{a}\right)}{4\pi|L|\left(E-\sqrt{E^{2}-m_{a}^{2}}\right)\sqrt{E^{2}-m_{a}^{2}}}$
×$\times e^{-\mathrm{i}(\sqrt{E^{2}-m_{a}^{2}}-g_{a})|L|},$
$I_{4}$ $=\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a+}^{2}e^{\mathrm{i}\mathbf{qL}}}{E+E_{a+}-\mathrm{i}0}=-\frac{\left(E-\sqrt{E^{2}-m_{a}^{2}}\right)\left(\sqrt{E^{2}-m_{a}^{2}}+g_{a}\right)}{4\pi|L|\sqrt{E^{2}-m_{a}^{2}}}$
×$\times e^{-\mathrm{i}(\sqrt{E^{2}-m_{a}^{2}}+g_{a})|L|},$ (B.10)
and
$J_{2}=$ $\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a+}^{2}A_{a+}^{2}(\bm{\sigma}\hat{q})e^{\mathrm{i}\mathbf{qL}}}{E-E_{a+}+\mathrm{i}0}\approx-(\bm{\sigma}\hat{L})\frac{m_{a}^{2}\left(\sqrt{E^{2}-m_{a}^{2}}+g_{a}\right)}{4\pi|L|\left(E+\sqrt{E^{2}-m_{a}^{2}}\right)\sqrt{E^{2}-m_{a}^{2}}}$
×$\times e^{\mathrm{i}(\sqrt{E^{2}-m_{a}^{2}}+g_{a})|L|}=(\bm{\sigma}\hat{L})I_{2},$
$J_{3}=$ $\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a-}^{2}A_{a-}^{2}(\bm{\sigma}\hat{q})e^{\mathrm{i}\mathbf{qL}}}{E+E_{a-}-\mathrm{i}0}\approx(\bm{\sigma}\hat{L})\frac{m_{a}^{2}\left(\sqrt{E^{2}-m_{a}^{2}}-g\right)}{4\pi|L|\left(E-\sqrt{E^{2}-m_{a}^{2}}\right)\sqrt{E^{2}-m_{a}^{2}}}$
×$\times e^{-\mathrm{i}(\sqrt{E^{2}-m_{a}^{2}}-g_{a})|L|}=-(\bm{\sigma}\hat{L})I_{3},$
$J_{4}=$ $\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\frac{\lambda_{a+}^{2}(\bm{\sigma}\hat{q})e^{\mathrm{i}\mathbf{qL}}}{E+E_{a+}-\mathrm{i}0}\approx(\bm{\sigma}\hat{L})\frac{\left(E-\sqrt{E^{2}-m_{a}^{2}}\right)\left(\sqrt{E^{2}-m_{a}^{2}}+g_{a}\right)}{4\pi|L|\sqrt{E^{2}-m_{a}^{2}}}$
×$\times e^{-\mathrm{i}(\sqrt{E^{2}-m_{a}^{2}}+g_{a})|L|}=-(\bm{\sigma}\hat{L})I_{4}.$ (B.11)
In Eq. (B.11), we show the limit of great $L$. One can see in Eqs. (B.6) and (B), that $|I_{1,3}|\gg|I_{2,4}|$ for ultrarelativistic particles with $E\gg m_{a}$.
According to Eq. (2.10), the matrix element is proportional to the Fourier transform of the propagator in Eq. (B.1) averaged over the lepton wavefunctions. For simplicity, we neglect the mixing matrix in Eq. (2.10). Considering ultrarelativistic left leptons in initial and final states with $\mathbf{p}\parallel\mathbf{L}$, as well as using Eqs. (B.6), (B.9), (B), and (B.11), one has that
$\mathcal{M}=$ $\kappa_{-}^{\dagger}(\mathbf{p})\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}S(E,\mathbf{q})e^{\mathrm{i}\mathbf{qL}}\kappa_{-}(\mathbf{p})$
$\approx\frac{\mathrm{i}}{2}\kappa_{-}^{\dagger}(\mathbf{p})\left\{(I_{1}+I_{2})[1-(\bm{\sigma}\hat{L})]+(I_{3}+I_{4})[1+(\bm{\sigma}\hat{L})]\right\}\kappa_{-}(\mathbf{p})$
$\approx\mathrm{i}I_{1}\approx-\frac{\mathrm{i}E}{4\pi|L|}e^{\mathrm{i}(\sqrt{E^{2}-m_{a}^{2}}-g_{a})|L|},$ (B.12)
since $[1+(\bm{\sigma}\hat{L})]\kappa_{-}(\mathbf{p})=0$ and $[1-(\bm{\sigma}\hat{L})]\kappa_{-}(\mathbf{p})=2\kappa_{-}(\mathbf{p})$.
Thus, if one averages the matrix element over the left-polarized states corresponding to ultrarelativistic leptons, the main contribution results from the $I_{1}$ and $J_{1}$ terms in the neutrino propagator in Eq. (A.10). The contribution of the propagator in Eq. (A.11) to the matrix element is suppressed by the factor $m_{a}/E$, which is small for ultrarelativistic neutrinos. These are the main results of this Appendix.